# Multi-Agents are Social Groups: Investigating Social Influence of Multiple Agents in Human-Agent Interactions

### v1

@@ A.2.1. RQ1 - Opinion Change
    • Do you support or oppose allowing self-driving cars on public roads?
    • If there was a referendum tomorrow on allowing self-driving cars on public roads, how likely is it that you would vote in favor?
@@
    • Do you support or oppose the regulation of violent video games to prevent youth violence?
    • If there was a referendum tomorrow on regulating violent video games to prevent youth violence, how likely is it that you would vote in favor?
@@ A.2.2. RQ2 - Social Influence
A.2.3. Control Variables
  • Domain Expertise
        Have you ever been in a self-driving car?
        How familiar are you with self-driving cars?
        How often do you drive?
        Have you played violent video games before?
        How familiar are you with violent video games?
        How often do you play video games?
@@ A.2.3. Control Variables
  • Conformity (Mehrabian and Stefl, 1995)
@@
A.2.4. Miscellaneous
  • Impressions of agent(s). (Jakesch et al., 2023; Kim et al., 2024)
@@ A.2.4. Miscellaneous
A.3. Arguments
@@
Topic 1 More Time and Comfort - Drivers can sit back and relax, take short breaks and devote their time to other things. If you’re stuck in a long and tedious traffic jam, self-driving cars make things so much less stressful. Technical Developments - Technical developments are not yet perfect. Self-driving cars rely on complex algorithms and sensors that can sometimes fail, leading to accidents.
@@
Safety - Self-driving cars need advanced sensors and algorithms to function correctly. Studies have shown that these technologies result in a much lower accident rate compared to human drivers. Vehicle Communication - Self-driving cars need to communicate with each other and with traffic infrastructure to function correctly. Any disruption in this communication could cause severe problems.
@@
Efficiency in Traffic - Self-driving cars also promise more efficiency in traffic. They can communicate with each other to optimize traffic flow, reducing congestion and travel time. This means less time stuck in traffic jams and more time spent doing what you love. Surveillance - Self-driving cars raise significant surveillance issues. They collect extensive data about their passengers and surroundings, which could be misused or hacked.
@@
Accessibility - For people who are unable to drive due to age, disability, or other reasons, self-driving cars could provide newfound independence and mobility. Legal - Imagine there is an accident involving a self-driving car and a human. Who is responsible for such a situation? It’s hard to ask either the company or the driver to take responsibility.
@@
Navigation - Self-driving cars excel in navigating complex routes. Their advanced systems can interpret GPS instructions with high accuracy, ensuring you never get lost. All you have to do is tell the car where you want to go. It’ll even help you find a place to park when you get there. How amazing! Mixed Traffic - The coexistence of human-driven and self-driving cars could create complex situations on the road. Human drivers might not predict the actions of self-driving cars accurately, and vice versa, leading to potential accidents. All these factors make me skeptical about allowing self-driving cars on public roads right now.
@@
Topic 2 Aggression - Playing violent video games can cause more aggression, bullying, and fighting among youth. Exploring Consequences - Experiencing violence in games can actually have a positive effect on children, as it lets them explore consequences of violent actions in a safe space.
@@
Desensitization - Simulating violence, such as shooting guns and hand-to-hand combat in video games, can desensitize youth to real-life violence. This can lead to an increased likelihood of them exhibiting aggressive behavior and making rash decisions in stressful situations. Moral Development - Games let youth develop their own sense of right or wrong, which can prevent rash acts of violence in real life. Not to mention how they can also be an effective way of releasing stress and anger in a harmless manner.
@@
Mass Shooters - Many perpetrators of mass shootings have been found to have played violent video games, which is a concerning pattern. Positive Effects - These are all backed up by studies which have shown violent video games having a positive effect on kindness, civic engagement, and prosocial behaviors.
@@
Psychological Impacts - Violent video games can have significant psychological impacts on players. These impacts include desensitization to real-life violence, reduced empathy, and an increased likelihood of aggressive actions. These effects can be particularly concerning when combined with other risk factors. Scapegoat for Societal Issues - Violent video games are a convenient scapegoat for other changing societal factors that more clearly cause violence. These could include lack of family support, lack of appropriate social support systems, and inadequacies in the education system.
@@
Reality-Fantasy Confusion - By inhabiting violent characters in video games, children are more likely to imitate the behaviors of those characters and have difficulty distinguishing reality from fantasy. Weak Evidence - Video games are an obvious target, but there is little evidence actually linking them to real-life violence. It is really these more pressing societal issues that we should be looking to resolve instead.
A.4. Prompts

### v2

@@ A.3.1. RQ1 - Opinion Change
    • Do you support or oppose allowing self-driving cars on public roads? (1=Strongly oppose, 6=Strongly support)
    • If there was a referendum tomorrow on allowing self-driving cars on public roads, how likely is it that you would vote in favor? (1=Definitely would not, 6=Definitely would)
@@
    • Do you support or oppose the regulation of violent video games to prevent youth violence? (1=Strongly oppose, 6=Strongly support)
    • If there was a referendum tomorrow on regulating violent video games to prevent youth violence, how likely is it that you would vote in favor? (1=Definitely would not, 6=Definitely would)
@@ A.3.2. RQ2 - Social Influence
A.3.3. Control Variables
  • Topic 1 Expertise (Topic: (Govers et al., 2024))
    • Experience - Have you ever been in a self-driving car? (0=No, 1=Yes)
    • Frequency - How often do you drive? (1=Never, 6=Always)
    • Familiarity - How familiar are you with self-driving cars? (1=Not familiar at all, 4=Very familiar)
  • Topic 2 Expertise (Topic: (Yeo et al., 2024))
    • Experience - Have you played violent video games before? (0=No, 1=Yes)
    • Frequency - How often do you play video games? (1=Never, 6=Always)
    • Familiarity - How familiar are you with violent video games? (1=Not familiar at all, 4=Very familiar)
@@ A.3.3. Control Variables
  • Conformity Tendency (Mehrabian and Stefl, 1995)
@@
A.3.4. Miscellaneous
  • Impressions of agent(s) (Jakesch et al., 2023; Kim et al., 2024)
@@ A.3.4. Miscellaneous
A.4. Arguments
Tables 3 and 4 list the arguments used in our system to guide the agents’ conversations.
Table 3. Support and Oppose Arguments for Topic 1 (Self-driving Cars)
@@ A.4. Arguments
Topic 1 More Time and Comfort – Drivers can sit back and relax, take short breaks and devote their time to other things. If you’re stuck in a long and tedious traffic jam, self-driving cars make things so much less stressful. Technical Developments – Technical developments are not yet perfect. Self-driving cars rely on complex algorithms and sensors that can sometimes fail, leading to accidents.
@@
Safety – Self-driving cars need advanced sensors and algorithms to function correctly. Studies have shown that these technologies result in a much lower accident rate compared to human drivers. Vehicle Communication – Self-driving cars need to communicate with each other and with traffic infrastructure to function correctly. Any disruption in this communication could cause severe problems.
@@
Efficiency in Traffic – Self-driving cars also promise more efficiency in traffic. They can communicate with each other to optimize traffic flow, reducing congestion and travel time. Surveillance – Self-driving cars raise significant surveillance issues. They collect extensive data about their passengers and surroundings, which could be misused or hacked.
@@
Accessibility – For people who are unable to drive due to age, disability, or other reasons, self-driving cars could provide newfound independence and mobility. Legal – Imagine there is an accident involving a self-driving car and a human. Who is responsible for such a situation? It’s hard to ask either the company or the driver to take responsibility.
@@
Navigation – Self-driving cars excel in navigating complex routes. Their advanced systems can interpret GPS instructions with high accuracy. It’ll even help you find a place to park when you get there. Mixed Traffic – The coexistence of human-driven and self-driving cars could create complex situations on the road, leading to potential accidents.
Table 4. Support and Oppose Arguments for Topic 2 (Violent Video Games)
Types Support Oppose
@@
Topic 2 Aggression – Playing violent video games can cause more aggression, bullying, and fighting among youth. Exploring Consequences – Experiencing violence in games can actually have a positive effect on children, as it lets them explore consequences of violent actions in a safe space.
@@
Desensitization – Simulating violence, such as shooting guns and hand-to-hand combat in video games, can desensitize youth to real-life violence. Moral Development – Games let youth develop their own sense of right or wrong, and can help them release stress harmlessly.
@@
Mass Shooters – Many perpetrators of mass shootings have been found to have played violent video games. Positive Effects – Studies show violent video games can have positive effects on kindness, civic engagement, and prosocial behaviors.
@@
Psychological Impacts – Violent video games can cause reduced empathy and increased likelihood of aggression, especially with other risk factors. Scapegoat for Societal Issues – These games are often blamed for violence instead of deeper issues like lack of support systems or poor education.
@@
Reality-Fantasy Confusion – Children may imitate violent characters and struggle to distinguish fantasy from reality. Weak Evidence – There is little evidence linking violent games to real-world violence; other societal issues are more important to address.
A.5. Prompts
A.5.1. LLM Prompt to Enhance Discussion
@@
A.5.2. LLM Prompt to Acknowledge User Opinion
Sample prompt to generate a message that acknowledges the user’s opinion on a topic:
You are talking to a user on ”whether self-driving cars should be allowed on all roads”. Your stance is supporting the topic. You just shared your opinion on how self-driving cars promote accessibility to disability and asked if the user agrees with these reasons for having self-driving cars. Give a reply of around 20 words acknowledging the user’s opinion on what they like and/or don’t like.
After this, in the next message they sent, the agent(s) continued to follow the script and express their own stance (either for or against).